You are given an unknown salt that contains a mixture of FeCl2 (s) and FeCl3 (s). Describe a procedure for determining the mass percentage of each compound salt
Answer #1

We know the total mass of the mixture containing FeCl2 and FeCl3. These two salt are differed only by oxidation state.

We can use precipitation titration. The titration involves formation of precipitation of one of the products.

Now we can allow to react both salt with AgNO3. Due to difference in oxidation state of Fe, both salt forms different numbers of moles of AgCl (s). Lets show the reactions of these with AgNO3.

FeCl2:

FeCl2 (aq) + 2 AgNO3 (aq) ---- > 2AgCl (s) + Fe(NO3)2 (aq)

FeCl3:

FeCl3 (aq) + 3 AgNO3 (aq) ---- > 3AgCl (s) + Fe(NO3)3 (aq)

In both the reactions above we get moles of AgCl(s)

Total mixture is now dissolved in water and we allow it to react with AgNO3. After completion of titration we find moles of AgCl(s)

According to above reaction we can generate following two mole relations

5 mol AgCl is equivalent to 2/5 moles of FeCl2

5 moles of AgCl is equivalent to 3/5 moles of FeCl3

In other words we can say that

5 moles of AgCl are produced from 2/5 moles of FeCl3 and 3/5 moles of FeCl3

Lets make its mathematical expression.

1 moles of AgCl = 2/5 FeCl2 +3/5 FeCl3

0.4 mol FeCl2 + 0.6 moles FeCl3 = 1 mol of AgCl

Or percent of FeCl2 required = 40 % and that of FeCl3 is 60 %

So if we get x moles of AgCl then

Moles of FeCl2 = x * 40 / 100 = 0.4 x moles of AgCl

And moles of FeCl3 = x * 60/100 = 0.6 x moles of AgCl

Once we get their moles then we can use molar mass and find their corresponding masses.

And mass percent of FeCl2 = (mass of FeCl2 / total mass) x 100

And mass percent of FeCl3 = (mass of FeCl3 / total mass) x 100